# NeuroAMP: A Novel End-to-end General Purpose Deep Neural Amplifier for Personalized Hearing Aids

### v1

@@ 4.3.3 Waveform Analysis
To further validate our model’s performance, we conducted an in-depth analysis encompassing waveform comparison, spectral analysis, and frequency band energy over time. We compared the unprocessed NAL-R, NAL-R+WDRC, and NeuroAMP amplified signals. Fig. 10 presents the waveform comparison, revealing that the Transformer model’s output closely aligns with the ground truth (NAL-R+WDRC) signal. The similarity in waveform patterns demonstrates the effectiveness of our model in mimicking the ground-truth waveform, as highlighted in the red boxes.
@@ 4.3.4 Spectrogram Analysis
Fig. 11 presents the spectral comparison for unprocessed NAL-R, NAL-R+WDRC, and NeuroAMP amplified signals. The spectral analysis aligns the findings from the waveform comparison, indicating that the Transformer model’s spectral characteristics closely resemble those of the NAL-R+WDRC amplified signal. This consistency across both waveform and spectral domains highlights NeuroAMP’s capability to replicate the performance characteristics of the ground-truth signals.
@@ 4.3.5 Frequency Band Analysis
Fig. 12 illustrates the comparative analysis of frequency band energy over time for unprocessed, NAL-R, NAL-R+WDRC, and NeuroAMP-processed signals. This analysis provides insights into how each processing method affects the time-frequency distribution, focusing on specific frequency bands and their energy distribution. In the low-frequency band (0-500 Hz), the unprocessed signal shows the lowest energy levels with significant fluctuations over time. The NAL-R processed signal exhibits higher energy levels, indicating effective amplification in this region. The NAL-R+WDRC processing maintains these elevated energy levels while adding dynamic range compression, which is visible as slight variations over time. Notably, the NeuroAMP-processed signal closely mirrors the energy pattern of the NAL-R+WDRC, demonstrating its ability to effectively replicate low-frequency enhancements.
In the mid-frequency band (500-2000 Hz), the unprocessed signal shows the lowest energy levels. The NAL-R processing significantly increases energy in this range, enhancing mid-frequency components crucial for speech intelligibility. While similar to NAL-R in energy levels, the NAL-R+WDRC processing introduces slight variations due to dynamic range compression. The NeuroAMP model closely aligns with the NAL-R+WDRC energy distribution, demonstrating a high degree of accuracy in replicating mid-frequency enhancements.
In the high-frequency band (2000-8000 Hz), the unprocessed signal maintains low energy levels. The NAL-R processed signal shows significant amplification, enhancing high-frequency components crucial for speech intelligibility. As with other bands, the NAL-R+WDRC processing introduces minor variations due to compression effects. The NeuroAMP-processed signal closely mirrors the NAL-R+WDRC pattern, effectively replicating both high-frequency amplification and compression.
@@ 4.3.6 Gain Function Analysis
In Panel (d), the NAL-R processed gain (blue line) shows a more pronounced increase, particularly at higher frequencies, to address the greater hearing loss shown in Panel (c). The NAL-R+WDRC processed gain (red line) continues to follow the NAL-R curve, with compression effects providing a stable gain distribution. The NeuroAMP-processed gain (green line with triangles) closely aligns with the NAL-R+WDRC gain pattern, demonstrating NeuroAMP’s effective replication of the gain characteristics. This analysis highlights that the NeuroAMP model successfully replicates the gain features of NAL-R+WDRC processing across both hearing loss profiles. Its ability to offer appropriate amplification and compression tailored to specific hearing loss profiles suggests it as a viable alternative for hearing aid signal processing. By matching the performance of established methods, the NeuroAMP model ensures comparable speech intelligibility and listening comfort for users, similar to traditional methods.
@@
4.4 Subjective Evaluations
This section presents a preference test comparing our proposed NeuroAMP model with the traditional NAL-R+WDRC processing. The evaluation involved 20 subjects, evenly split between males and females, all confirmed to have normal hearing through audiometric testing. To simulate hearing loss, we used the MSBG hearing loss simulator [59], enabling a realistic assessment of processed audio as perceived by individuals with hearing loss. Each subject listened to pairs of audio samples processed by the NAL-R+WDRC and NeuroAMP models, followed by the MSBG hearing loss simulator. They were asked to determine whether sample A was better, sample B was better, or if the samples were similar and indistinguishable. The sample order was randomized, and the tests were conducted in a double-blind manner. We selected a diverse range of samples from our test dataset to ensure a comprehensive evaluation, including 10 clean speech samples, 10 noisy speech samples, 10 speech samples enhanced by the FCN model, 10 speech samples enhanced by the LSTM model, 5 clean music samples, and 5 noisy music samples. Participants evaluated each pair based on intelligibility (clarity of speech) and quality (overall audio quality, including naturalness and pleasantness).
@@ 4.4 Subjective Evaluations
Fig. 14 shows the total number of evaluation results, aggregating preferences and ratings across different sample types. Detailed visual representations of the evaluation results for each sample type—clean speech, noisy speech, enhanced speech by FCN, enhanced speech by LSTM, clean music, and noisy music—are shown in Fig. 15. From Fig. 14, we observe that ”Similar” responses significantly outnumber those for ”NAL-R+WDRC” and ”NeuroAMP.” Fig. 15 shows a consistent trend: ”Similar” decisions prevail across all conditions, including speech, noisy speech, enhanced speech, and music scenarios. These results confirm that the proposed NeuroAMP model effectively simulates the performance of NAL-R+WDRC and demonstrates its capability to deploy a personalized deep learning-based amplification approach.
4.5 Denoising NeuroAMP Performance Evaluation
This section evaluates the performance of the Denoising NeuroAMP model, which combines speech enhancement and personalized amplification in a single unified process. The evaluation focuses on the VoiceBank+DEMAND dataset, a widely used benchmark for speech enhancement tasks. We compare Denoising NeuroAMP to two baselines:
  • Denoising>NAL-R+WDRC: This is a two-stage baseline that first uses a separate LSTM-based speech enhancement model to enhance noisy speech, and then uses NAL-R+WDRC for amplification.
We use HASPI and HASQI as evaluation metrics. These metrics are commonly used to assess the intelligibility and quality of processed speech for hearing aid users. Table VI lists the HASPI and HASQI scores of Denoising NeuroAMP and the two baselines on the noisy test set of the VoiceBank+DEMAND dataset.
Table 6: Performance Comparison of Denoising NeuroAMP on the VCTK Noisy Test Set
Method HASPI HASQI
NAL-R+WDRC (Noisy Input) 0.89 0.55
Denoising>NAL-R+WDRC (Two-Stage) 0.85 0.30
Denoising NeuroAMP (End-to-End) 0.90 0.59
The results show that Denoising NeuroAMP outperforms both baselines in both HASPI and HASQI scores. Specifically, Denoising NeuroAMP achieves a HASPI score of 0.90 and a HASQI score of 0.59, while NAL-R+WDRC has scores of 0.89 and 0.55, and the two-stage Denoising>NAL-R+WDRC baseline has scores of 0.85 and 0.30. The improvements of Denoising NeuroAMP over Denoising>NAL-R+WDRC (a standard two-stage approach commonly used in hearing aids) are confirmed by a t-test to be statistically significant. These findings suggest that Denoising NeuroAMP, an integrated approach that jointly optimizes speech enhancement and personalized amplification, is more effective than applying these processes separately. The lower performance of the two-stage baseline, particularly with respect to HASQI, highlights the potential drawbacks of a disjointed approach, where errors in the speech enhancement stage may negatively impact the subsequent amplification stage. Denoising NeuroAMP’s superior performance confirms its potential to improve speech intelligibility and quality for hearing aid users in noisy environments. By learning a unified denoising and amplification process tailored to an individual’s hearing profile, Denoising NeuroAMP provides an alternative solution to more effective hearing aid technology.
@@ 5 Conclusion
This research introduces NeuroAMP, a novel deep learning-based system for personalized hearing aid amplification, and Denoising NeuroAMP, an extension that integrates noise reduction for better performance in adverse environments. We investigated four neural network architectures: CNN, LSTM, CRNN, and Transformer, using spectral features and audiograms as inputs. Extensive evaluation using HASPI, HASQI, and HAAQI, together with statistical measures like LCC, SRCC, and MSE, demonstrated the superior performance of the Transformer within the NeuroAMP framework. Notably, NeuroAMP achieved SRCC scores of 0.9810 (HASQI) and 0.9914 (HASPI) on the unseen VCTK dataset, and 0.9892 (HAAQI) on the MUSDB18-HQ dataset, validating the effectiveness of our data augmentation strategy. Next, Subjective listening tests confirmed that both NeuroAMP produce outputs that are perceptually similar to those of traditional methods while offering better adaptability. Furthermore, Denoising NeuroAMP outperformed conventional NAL-R+WDRC and two-stage baselines on the VoiceBank+DEMAND dataset, highlighting its potential for enhanced speech intelligibility in noisy environments. In conclusion, this work demonstrates the potential of NeuroAMP and Denoising NeuroAMP to perform a personalized, user-centric approach to amplification and noise reductions. In our future work, we will focus on refining these models, exploring real-time implementation, incorporating user feedback, and integrating other essential modules, such as a hearing loss model for compensation, to achieve a better personalized hearing assistance solution.

### v2

@@ 4.3.5 Waveform Analysis
To further validate our model’s performance, we conducted an in-depth analysis encompassing waveform comparison, spectral analysis, and frequency band energy over time. Using a moderate-severity sloping audiogram with high-frequency hearing loss as the listener profile, we compared the unprocessed NAL-R, NAL-R+WDRC, and NeuroAMP amplified signals. Fig. 12 presents the waveform comparison, revealing that the Transformer model’s output closely aligns with the ground-truth (NAL-R+WDRC) signal. The similarity in waveform patterns demonstrates our model’s effectiveness in accurately replicating the ground-truth waveform, as indicated by the highlighted regions (red boxes).
@@ 4.3.6 Spectrogram Analysis
Fig. 13 presents the spectral comparison for unprocessed NAL-R, NAL-R+WDRC, and NeuroAMP amplified signals. The spectral analysis aligns with the findings from the waveform comparison, indicating that the Transformer model’s spectral characteristics closely resemble those of the NAL-R+WDRC amplified signal. This consistency across both waveform and spectral domains highlights NeuroAMP’s capability to replicate the performance characteristics of the ground-truth signals.
@@ 4.3.7 Frequency Band Analysis
Fig. 14 illustrates the comparative analysis of frequency band energy over time for unprocessed, NAL-R, NAL-R+WDRC, and NeuroAMP-processed signals. This analysis provides insights into how each processing method affects the time-frequency distribution, focusing on specific frequency bands and their energy distribution. In the low-frequency band (0–500 Hz), all signals show similar overall energy trends. NAL-R and NAL-R+WDRC maintain comparable low-frequency energy, with the latter adding subtle compression-induced smoothing. Notably, the NeuroAMP-processed signal closely mirrors the energy pattern of the NAL-R+WDRC, demonstrating its ability to effectively replicate low-frequency signal processing.
In the mid-frequency band (500–2000 Hz), the unprocessed signal shows the lowest energy levels. The NAL-R processing significantly increases energy in this range, enhancing mid-frequency components crucial for speech intelligibility. While similar to NAL-R in energy levels, the NAL-R+WDRC processing introduces slight variations due to dynamic range compression. The NeuroAMP model closely aligns with the NAL-R+WDRC energy distribution, demonstrating high accuracy in replicating mid-frequency amplification and compression.
In the high-frequency band (2000-8000 Hz), the unprocessed signal maintains low energy levels. The NAL-R processed signal shows significant amplification, amplifying high-frequency components crucial for speech intelligibility. As with other bands, NAL-R+WDRC processing introduces minor variations due to compression effects. The NeuroAMP-processed signal closely mirrors the NAL-R+WDRC pattern, effectively replicating both high-frequency amplification and compression.
@@ 4.3.8 Gain Function Analysis
In panel (d), the NAL-R processed gain (blue line) shows a more pronounced increase, particularly at higher frequencies, to address the greater hearing loss shown in Panel (c). The NAL-R+WDRC processed gain (red line) continues to follow the NAL-R curve, with compression effects providing a stable gain distribution. The NeuroAMP-processed gain (green line with triangles) closely aligns with the NAL-R+WDRC gain pattern, demonstrating NeuroAMP’s effective replication of the gain characteristics. This analysis highlights that the NeuroAMP model successfully replicates the gain features of NAL-R+WDRC processing across both hearing loss profiles. Its ability to provide appropriate amplification and compression tailored to specific hearing loss profiles indicates that it is a viable alternative for hearing aid signal processing. By matching the performance of established methods, the NeuroAMP model ensures comparable speech intelligibility and listening comfort for users, similar to traditional methods.
4.4 Denoising NeuroAMP Performance Evaluation
This section evaluates the performance of the Denoising NeuroAMP model, which combines speech enhancement and personalized amplification in a single unified model. The evaluation was conducted on the VoiceBank-DEMAND dataset. We compare Denoising NeuroAMP to two baselines:
[…]
Table VI lists the HASPI and HASQI scores of Denoising NeuroAMP and the two baselines on the noisy test set of the VoiceBank-DEMAND dataset.
Table 6: Performance Comparison of Denoising NeuroAMP on the VoiceBank-DEMAND Noisy Test Set
Method HASPI HASQI
NAL-R+WDRC (Noisy Input) 0.89 0.55
[…]
Denoising NeuroAMP (End-to-End) 0.90 0.59
[…]
4.5 Subjective Evaluations
4.5.1 Perceptual Similarity Analysis
This section presents a preference test comparing our proposed NeuroAMP model with the traditional NAL-R+WDRC processing. The evaluation involved 20 subjects, evenly split between males and females, all confirmed to have normal hearing through audiometric testing  [^2] All ethical and experimental procedures were approved by the Institutional Review Board (IRB) on Biomedical Science Research at Academia Sinica under protocol number AS-IRB-BM-24026. All experiments were conducted in accordance with the approved IRB protocols. To simulate hearing loss, we used the MSBG hearing loss simulator [58], enabling a realistic assessment of processed audio as perceived by individuals with hearing loss. Each subject listened to pairs of audio samples processed by either the NAL-R+WDRC or NeuroAMP models, which were then further processed using the MSBG hearing loss simulator. They were asked to determine whether sample A was better, sample B was better, or if the samples were similar and indistinguishable. The sample order was randomized, and the tests were conducted in a double-blind manner. We selected a diverse set of audio samples from our test dataset to ensure a comprehensive evaluation, including 10 clean speech samples, 10 noisy speech samples, 10 speech samples enhanced by the FCN model, 10 speech samples enhanced by the LSTM model, 5 clean music samples, and 5 noisy music samples. Participants evaluated each pair based on intelligibility (clarity of speech) and quality (overall audio quality, including naturalness and pleasantness).
@@ 4.5.1 Perceptual Similarity Analysis
Fig. 16 shows the evaluation results, aggregating preferences and ratings across different sample types. Proportions for each response category are shown with 95% confidence intervals, and statistical significance was assessed using a chi-square goodness-of-fit test, with the resulting p-value annotated in the figure. Detailed visual representations of the evaluation results for each sample type—clean speech, noisy speech, enhanced speech by FCN, enhanced speech by LSTM, clean music, and noisy music—are shown in Fig. 17. As shown in Fig. 16, responses indicating ’Similar’ significantly outnumbered preferences for either ’NAL-R+WDRC’ or ’NeuroAMP’. Fig. 17 shows a consistent trend: ”Similar” decisions prevail across all conditions, including speech, noisy speech, enhanced speech, and music scenarios. These results confirm that the proposed NeuroAMP model effectively replicates the perceptual characteristics of NAL-R+WDRC, demonstrating the viability of a deep learning-based amplification approach.
4.5.2 Mean Opinion Score (MOS)
To quantitatively assess the perceived audio quality of the processed samples, we conducted a MOS evaluation involving 30 listeners. The audio samples were processed using three amplification approaches: the traditional NAL-R+WDRC processing, the proposed NeuroAMP model, and Denoising NeuroAMP. Subsequently, each audio sample was processed using the MSBG hearing loss simulator to emulate auditory experiences representative of individuals with hearing impairment. The listeners were asked to evaluate audio samples from three distinct categories: clean speech, noisy speech, and music. Each participant rated perceived audio quality using a MOS scale from 1 (”poor” quality) to 5 (”excellent” quality). The evaluation results, including average scores and 95% confidence intervals, are presented in Fig. 18. Statistical significance was assessed using one-way ANOVA and pairwise t-tests, with significant differences indicated by asterisks (*) above the bars.
Figure 18: MOS comparison for perceived audio quality across three audio categories (clean speech, noisy speech, and music) processed using NAL-R+WDRC, NeuroAMP, and Denoising NeuroAMP methods. Bars represent mean scores with 95% confidence intervals. Significant differences between processing methods (p << 0.05) are marked by asterisks (*). The hearing-impaired speech signals were simulated using the MSBG model.
As shown in Fig. 18, the three processing approaches demonstrate comparable performance for clean speech, with overlapping confidence intervals. Specifically, Denoising NeuroAMP achieved the highest MOS (3.71), closely followed by NeuroAMP (3.64) and the conventional NAL-R+WDRC method (3.62). In noisy speech conditions, Denoising NeuroAMP achieves the best performance with an average MOS of 3.57, and the pairwise comparison shows this difference to be statistically significant compared to the other methods (as indicated by asterisks). NeuroAMP and NAL-R+WDRC also show comparable performance, with NeuroAMP achieving a higher MOS than NAL-R+WDRC (2.90 vs. 2.82). For music, NeuroAMP and NAL-R+WDRC produced closely matched results, with NeuroAMP slightly outperforming NAL-R+WDRC (2.80 vs. 2.77).
Overall, the MOS evaluation confirms that the proposed NeuroAMP model and the conventional NAL-R+WDRC processing offer similar perceptual quality across various audio categories. Notably, the addition of denoising capabilities through Denoising NeuroAMP significantly improves audio quality in noisy speech scenarios, highlighting the potential to further enhance deep learning-based amplification frameworks.
Finally, we recruited 7 individuals with hearing loss to participate in the MOS listening test. Each subject had a specific audiogram, and the evaluation employed the same set of speech samples as used in Fig.18. Prior to the listening tests, audiograms for each participant’s target ear were obtained. Based on these audiograms, the speech signals were processed using NAL-R+WDRC, NeuroAMP, and Denoising NeuroAMP, and then presented to the participants without the use of hearing aids. The resulting MOS scores (mean ± 95% confidence intervals) are shown in Fig. 19. Statistical significance was assessed using one-way ANOVA and Bonferroni-corrected pairwise t-tests, with significant differences marked by asterisks (*) in the figure. The trends observed in Fig. 19 align with those in Fig. 18. In clean and music conditions, NAL-R+WDRC and NeuroAMP demonstrated comparable performance, with NeuroAMP slightly outperforming NAL-R+WDRC. Under noisy conditions, Denoising NeuroAMP exhibited significant improvements over the other two methods.
Figure 19: MOS comparison for perceived audio quality across three audio categories (clean speech, noisy speech, and music) processed using NAL-R+WDRC, NeuroAMP, and Denoising NeuroAMP methods. Seven subjects with hearing loss participated in the listening test.
@@ 5 Conclusion
This research introduces NeuroAMP, a novel deep learning-based system for personalized hearing aid amplification, and Denoising NeuroAMP, an extension that integrates noise reduction for better performance in adverse environments. We investigated four neural network architectures: CNN, LSTM, CRNN, and Transformer, using spectral features and audiograms as inputs. Extensive evaluation using HASPI, HASQI, and HAAQI, together with statistical measures like LCC, SRCC, and MSE, demonstrated the superior performance of the Transformer within the NeuroAMP framework. Notably, NeuroAMP achieves SRCC scores of 0.9810 (HASQI) and 0.9914 (HASPI) on the unseen VoiceBank-DEMAND dataset, and 0.9892 (HAAQI) on the MUSDB18-HQ dataset, validating the effectiveness of our data augmentation strategy. Subjective listening tests further confirmed that NeuroAMP produces outputs perceptually similar to traditional methods while providing greater adaptability. Furthermore, Denoising NeuroAMP outperformed conventional NAL-R+WDRC and two-stage baselines on the VoiceBank-DEMAND dataset, highlighting its potential for enhanced speech intelligibility in noisy environments. In conclusion, this work demonstrates the potential of NeuroAMP and Denoising NeuroAMP as personalized, user-centric solutions for amplification and noise reduction. In our future work, we will focus on refining these models, exploring real-time implementation, incorporating user feedback, and integrating other essential modules, such as a hearing loss model for compensation, to achieve improved personalized hearing assistance solutions.